OpenStudy (anonymous):

zeros of x^2+5x-43=5x-6

OpenStudy (anonymous):

We know:\[ x^2+5x-43=5x-6\implies\\ x^2-43+6=x^2-37=0\implies\\ x=\pm \sqrt{37} \]Thus, we are done.